### 01 Feb 2024 v1

@@ 1 Introduction
Uncertainty quantification plays a crucial role in time series forecasting, particularly for downstream applications that require more nuanced insights than simple point forecasts to guide decision making. Probabilistic time series forecasting based on deep learning (DL) has gained substantial attention in recent years, due to its outstanding performance in capturing complex and nonlinear dependencies in time series data and its ability to provide the probability distribution of target variables Benidis et al. (2022). For multivariate time series, recent neural network models (Salinas et al., 2019; Rasul et al., 2020; 2021; Drouin et al., 2022; Ashok et al., 2023) are typically framed as a regression problem $\mathbf{z}_{t}=f\left(\mathbf{x}_{t}\right)+\boldsymbol{\eta}_{t}$. In a simple scenario, the error term $\boldsymbol{\eta}_{t}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$, where $\boldsymbol{\Sigma}_{t}$ is the contemporaneous covariance matrix. A common assumption in these models is that errors are independent over time, which means $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any times $s$ and $t$ such that $s\neq t$. However, real-world data rarely conforms to this assumption—errors often exhibit a certain degree of autocorrelation due to factors such as the omission of essential covariates.
Addressing error autocorrelation is an important research question in statistical time series models. A common strategy is to assume that the error series is a temporal process, such as following an autoregressive integrated moving average (ARIMA) model (Hyndman & Athanasopoulos, 2018). The similar challenge also arises in deep learning models. Previous studies have sought to integrate autocorrelated errors into the training process through modifying the loss function (Sun et al., 2021; Zheng et al., 2023b). However, the aforementioned methods, which are based on deterministic output, are not readily applicable to probabilistic forecasting models, particularly in the multivariate setting. A notable exception to address error autocorrelation in univariate time series is the innovative batch training method proposed by Zheng et al. (2023a), which learns a dynamic covariance matrix of time-assorted errors in a mini-batch. The covariance matrix can then be used to calibrate the predictive distribution of time series variables. It is shown that the method can consistently enhance the performance of base probabilistic models when compared to the naive training approach (i.e., without considering autocorrelated errors). However, the approach of Zheng et al. (2023a) is only applicable to univariate probabilistic time series forecasting models (e.g., DeepAR (Salinas et al., 2020)).
@@ 2 Probabilistic Time Series Forecasting
Denote by $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$ the vector containing the values of multivariate time series at time step $t$, where $N$ is the number of time series. Given the observed history $\{\mathbf{z}_{t}\}_{t=1}^{T}$, probabilistic time series forecasting can be formulated as estimating the joint conditional distribution $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$ and $\mathbf{x}_{t}$ are some known time-dependent covariates (e.g., time of day and day of week). In other words, we are interested in forecasting $Q$ steps of future values given $P$ steps of historical values and covariates. The model can be further factorized as
@@
which becomes an autoregressive model that can be used for both one-step-ahead ($Q=1$) and multistep-ahead forecasting in a rolling manner. When used in multistep-ahead prediction, samples are drawn for the prediction range ($t\geq T+1$) and fed back for the next time step until the end of the prediction range. In neural networks, the conditioning is encapsulated into a state vector $\mathbf{h}_{t}$ of a transition dynamics $f_{\Theta}$ that evolves over time $\mathbf{h}_{t}=f_{\Theta}\left(\mathbf{h}_{t-1},\mathbf{z}_{t-1},\mathbf{x}_{t}\right)$. Therefore, Eq. (1) can be simplified as
@@
where $\mathbf{h}_{t}$ is mapped to the parameters of a chosen parametric distribution (e.g., multivariate Gaussian).
Existing probabilistic multivariate time series forecasting models assume that the error at each time step is independent, and thus the time series vector $\mathbf{z}_{t}$ follows a multivariate Gaussian distribution:
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ map the state vector $\mathbf{h}_{t}$ to the mean and covariance of a multivariate Gaussian distribution. This formulation is equivalent to decomposing $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$ with $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. The temporally independent error assumption corresponds to $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any time points $s$ and $t$ such that $s\neq t$. Fig. 1 provides an empirical example of the contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$. The results are calculated based on the prediction residuals of GPVar on the $\mathtt{m4\_hourly}$ dataset. While multivariate models primarily focus on modeling the contemporaneous covariance, it is evident that the residuals are not independent across time, as $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})\neq\mathbf{0}$.
@@ 3.1 Probabilistic Time Series Forecasting
Probabilistic forecasting seeks to model the probability distribution for the target variable(s) rather than producing only a point estimate as seen in deterministic forecasting. In general, there are two primary approaches to achieve this objective, either through parametric probability density functions (PDF) or through the quantile function (Benidis et al., 2022). One example is MQ-RNN (Wen et al., 2017), which directly generates quantile forecasts through a sequence-to-sequence (Seq2Seq) recurrent neural network (RNN) architecture. PDF-based approaches, which are widely adopted, assume a specific distribution form (e.g., Gaussian, Poisson) for the target variables and leverage neural networks to generate the distribution parameters. Notably, DeepAR (Salinas et al., 2020) utilizes an RNN architecture to model hidden state transitions. The multivariate version of DeepAR—GPVar (Salinas et al., 2019)—utilizes a Gaussian copula to transform the original observations into Gaussian random variables. Subsequently, the model assumes a joint multivariate Gaussian distribution on the transformed data.
In addition to outputting distribution parameters, neural networks can be used to generate probabilistic model parameters, as exemplified by the deep state space model (SSM) (Rangapuram et al., 2018), which uses RNN to generate SSM parameters for predicting samples. The normalizing Kalman filter (NKF) (de Bézenac et al., 2020) combines normalizing flows (NFs) with the classical linear Gaussian state space model (LGM) to model nonlinear dynamics and evaluate the probability density function of observations. In NKF, RNN is used to produce parameters for the LGM at each time step. Subsequently, the output of the LGM is transformed into the observations using NFs. The deep factor model (Wang et al., 2019) consists of a deterministic global component and a random component. The global component is parameterized by RNN for generating the global factors of time series, while the random component is any classical probabilistic time series model (e.g., Gaussian white noise process) for representing random effects.
@@ 3.2 Modeling Correlated Errors
Error correlation in time series is an extensively explored subject in the fields of econometrics and statistics (Prado et al., 2021; Hyndman & Athanasopoulos, 2018; Hamilton, 2020). Autocorrelation, also known as serial correlation, captures the temporal correlation of errors in a time series. On the other hand, contemporaneous correlation reflects the correlation among variables within the same time period. Classical statistical frameworks, including autoregressive (AR) and moving average (MA) models, along with their multivariate counterparts, have been firmly established to address these correlation challenges. Recent developments in DL-based time series models have also sought to explicitly model error correlation structures to enhance forecasting performance.
@@ 4 Our Method
Gaussian likelihood-based models are built on the assumption that $\boldsymbol{\eta}_{t}$ are independently distributed according to a multivariate Gaussian distribution. Model training can be achieved by maximizing the log-likelihood
$\mathcal{L}=\sum_{t=1}^{T}\log p\left(\mathbf{z}_{t}\mid\theta\left(\mathbf{h}_{t}\right)\right)$ (6)
$\propto\sum_{t=1}^{T}-\frac{1}{2}[\ln\lvert\boldsymbol{\Sigma}_{t}\rvert+(\mathbf{z}_{t}-\boldsymbol{\mu}_{t})^{\top}\boldsymbol{\Sigma}_{t}^{-1}(\mathbf{z}_{t}-\boldsymbol{\mu}_{t})].$
@@ 4.1 Training with Autocorrelated Errors
Building upon the concept of constructing mini-batches to address autocorrelated errors Zheng et al. (2023a), we extend this approach to the multivariate scenario. In many existing deep probabilistic time series models, including GPVar (Salinas et al., 2019), each training batch encompasses a set of time series slices covering a temporal length of $P+Q$, where $P$ represents the conditioning range, and $Q$ denotes the prediction range. Gaussian likelihood is assessed individually at each time step within the prediction range. However, this straightforward method fails to capture the serial correlation of errors among consecutive time steps. To address this limitation, we organize $D$ consecutive batches into a mini-batch, with each batch covering a temporal length of $P+1$ (i.e., $Q=1$). In simpler terms, a traditional training batch comprises data $\mathbf{z}_{{T-P+1}:{T+Q}}$, while a mini-batch consists of data $\{\mathbf{z}_{{T-P+t}:{T+t}}\}_{t=1}^{D}$, where $T$ denotes a randomly sampled time step that separates the conditioning range and prediction range. It is evident that the mini-batch serves as a reconstruction of the conventional training batch, covering the same time horizon if $D=Q$. An example of the collection of target time series vectors in a mini-batch of size $D$ (the time horizon we use to capture serial correlation) is given by:
@@
4.2 Multi-step Rolling Prediction
Autoregressive models perform multistep-ahead forecasting in an iterative way, wherein a sample is drawn at each time step and serves as input for the subsequent step until the desired prediction range is reached. Our approach enhances this process in a similar way as in Zheng et al. (2023a) by offering additional calibration based on the learned correlation matrix $\boldsymbol{C}_{t}$. Assuming observations are available up to time step $t$, and recalling that the errors in a mini-batch collectively follow a multivariate Gaussian distribution, we can derive the conditional distribution of $\boldsymbol{\eta}_{t+1}$ given errors in the past $(D-1)$ steps:
@@ 4.2 Multi-step Rolling Prediction
where $\boldsymbol{\eta}_{\text{obs}}=\operatorname{vec}\left(\left[\boldsymbol{\eta}_{t-D+2},\ldots,\boldsymbol{\eta}_{t-1},\boldsymbol{\eta}_{t}\right]\right)\in\mathbb{R}^{(D-1)B}$ is the collection of observed residuals, which will be available at forecasting step $t+1$. Here, $\boldsymbol{\Sigma}_{\text{obs}}$ corresponds to the $(D-1)B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that describes the covariance of $\boldsymbol{\eta}_{\text{obs}}$, and $\boldsymbol{\Sigma}_{*}$ corresponds to the $B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that describes the covariance of $\boldsymbol{\eta}_{t+1}$ and $\boldsymbol{\eta}_{\text{obs}}$, i.e., $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}=\begin{bmatrix}\boldsymbol{\Sigma}_{\text{obs}}&\boldsymbol{\Sigma}_{*}^{\top}\\ \boldsymbol{\Sigma}_{*}&\boldsymbol{\Sigma}_{t+1}\end{bmatrix}$. Note that for brevity we remove the time index $t$ in $\boldsymbol{\Sigma}_{\text{obs}}$, $\boldsymbol{\Sigma}_{*}$ and $\boldsymbol{\eta}_{\text{obs}}$. As $\boldsymbol{\mu}_{t+1}$ is a deterministic output from the base model, we can first draw a sample $\tilde{\boldsymbol{\eta}}_{t+1}$ from Eq. (13), and then transform it to the target variables by
@@
As can be seen, the final distribution for ${\mathbf{z}}_{t+1}$ becomes
@@
Following this approach, multistep-ahead forecasting can be achieved by treating the sample $\tilde{\boldsymbol{\eta}}_{t+1}$ as an observation and continuing the process to obtain a trajectory of $\{\tilde{\boldsymbol{\eta}}_{t+q}\}_{q=1}^{Q}$.

### 07 Feb 2024 v2

@@ 1 Introduction
Uncertainty quantification plays a crucial role in time series forecasting, particularly for downstream applications that require more nuanced insights than simple point forecasts to guide decision making. Probabilistic time series forecasting based on deep learning (DL) has gained substantial attention in recent years, due to its outstanding performance in capturing complex and nonlinear dependencies in time series data and its ability to provide the probability distribution of target variables (Benidis et al., 2022). For multivariate time series, recent neural network models (Salinas et al., 2019; Rasul et al., 2020, 2021; Drouin et al., 2022; Ashok et al., 2023) are typically framed as a regression problem $\mathbf{z}_{t}=f\left(\mathbf{x}_{t}\right)+\boldsymbol{\eta}_{t}$. In a simple scenario, the error term $\boldsymbol{\eta}_{t}$ follows a Gaussian distribution $\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$, where $\boldsymbol{\Sigma}_{t}$ is the contemporaneous covariance matrix. A common assumption in these models is that errors are independent over time, which means $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any times $s$ and $t$ such that $s\neq t$. However, real-world data rarely conforms to this assumption—errors often exhibit a certain degree of autocorrelation due to factors such as the omission of essential covariates.
Addressing error autocorrelation is a crucial aspect of research in classical time series models. A common strategy in this regard is dynamic regression, which assumes that the error series follows a temporal process, such as an autoregressive integrated moving average (ARIMA) model (Hyndman & Athanasopoulos, 2018). The similar challenge also arises in deep learning models. Previous studies have sought to integrate autocorrelated errors into the training process through modifying the loss function (Sun et al., 2021; Zheng et al., 2023b). However, the aforementioned methods, which are based on deterministic output, are not readily applicable to probabilistic forecasting models, particularly in the multivariate setting. A notable exception to address error autocorrelation in univariate time series is the innovative batch training method proposed by Zheng et al. (2023a), which learns a dynamic covariance matrix of time-assorted errors in a mini-batch. The covariance matrix can then be used to calibrate the predictive distribution of time series variables. It is shown that the method can consistently enhance the performance of base probabilistic models when compared to the naive training approach (i.e., without considering autocorrelated errors). However, the approach of Zheng et al. (2023a) is only applicable to univariate probabilistic time series forecasting models (e.g., DeepAR (Salinas et al., 2020)).
@@ 2 Probabilistic Time Series Forecasting
Denote $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$ as the vector containing the time series variables at time step $t$, where $N$ is the number of time series. Probabilistic time series forecasting can be formulated as estimating the joint conditional distribution $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$ based on the observed history $\{\mathbf{z}_{t}\}_{t=1}^{T}$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$ and $\mathbf{x}_{t}$ is the vector containing known covariates (e.g., time information and time series identifiers) for all time steps. In essence, the problem can be described as predicting the time series values for $Q$ future time steps, using all available covariates and $P$ steps of historical time series values:
@@
which results in an autoregressive model that is capable of multi-step forecasting in a rolling manner. When performing multi-step forecasting, samples will be drawn in the prediction range ($t\geq T+1$) and fed back for the next time step until the end of the desired prediction range. In neural networks, the conditioning information is typically integrated into a hidden state vector $\mathbf{h}_{t}$ governed by a transition dynamics $f_{\Theta}$ evolving over time $\mathbf{h}_{t}=f_{\Theta}\left(\mathbf{h}_{t-1},\mathbf{z}_{t-1},\mathbf{x}_{t}\right)$. Hence, Eq. (1) can be expressed more concisely:
@@
where $\mathbf{h}_{t}$ is used for generating the parameters of a parametric distribution (e.g., multivariate Gaussian).
Existing models usually assume that the error at each time step is independent, and thus the time series vector $\mathbf{z}_{t}$ follows a multivariate Gaussian distribution:
@@
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ map the hidden state $\mathbf{h}_{t}$ to the mean and covariance of a multivariate Gaussian distribution. This formulation is equivalent to decomposing $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$ with $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. The temporally independent error assumption corresponds to $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any time points $s$ and $t$ such that $s\neq t$. Fig. 1 provides an empirical example of the contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$. The results are calculated based on the prediction residuals of GPVar on the $\mathtt{m4\_hourly}$ dataset. While multivariate models primarily focus on modeling the contemporaneous covariance, it is evident that the residuals are not independent across time, as $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})\neq\mathbf{0}$.
@@ 3.1 Probabilistic Time Series Forecasting
Probabilistic forecasting quantifies uncertainty by characterizing the probability distribution for the target variable(s). There are two primary approaches to achieve this objective. One approach involves using the quantile function. For example, the MQ-RNN (Wen et al., 2017) directly generates quantile forecasts through a sequence-to-sequence (Seq2Seq) recurrent neural network (RNN) architecture. Another approach is based on probability density functions (PDF) of the target variables. PDF-based approaches assume a specific probabilistic models (e.g., Gaussian processes) for the target variables and utilize neural networks to generate the parameters of these probabilistic models. Notably, DeepAR (Salinas et al., 2020) utilizes an RNN architecture to model hidden state transitions. The multivariate version of DeepAR—GPVar (Salinas et al., 2019)—utilizes a Gaussian copula to transform the original observations into Gaussian random variables. Subsequently, the model assumes a joint multivariate Gaussian distribution on the transformed data.
In addition to outputting distribution parameters, neural networks can be used to generate probabilistic model parameters, as exemplified by the deep state space model (SSM) (Rangapuram et al., 2018), which uses RNN to generate SSM parameters for predicting samples. The Normalizing Kalman Filter (NKF), introduced by (de Bézenac et al., 2020), integrates normalizing flows (NFs) with the classical Linear Gaussian State Space Model (LGM) to capture nonlinear dynamics and assess the probability density function of observations. In NKF, RNN is used to produce parameters for the LGM at each time step. Subsequently, the output of the LGM is transformed into the observations using NFs. Wang et al. (2019) proposed the deep factor model that is comporised of a deterministic global component and a random component. The global component is parameterized by RNN for generating the global factors of time series, while the random component is any classical probabilistic time series model (e.g., Gaussian white noise process) for representing random effects.
@@ 3.2 Modeling Correlated Errors
Error correlation in time series is an extensively explored subject in the fields of econometrics and statistics (Prado et al., 2021; Hyndman & Athanasopoulos, 2018; Hamilton, 2020). Autocorrelation, or serial correlation, and contemporaneous correlation are two primary types of correlation. Autocorrelation refers to the temporal correlation of errors in a time series, while contemporaneous correlation captures the correlation among variables within the same time period. Classical statistical frameworks, including autoregressive (AR) and moving average (MA) models, along with their multivariate counterparts, have been firmly established to address these correlation challenges. Recent developments in DL-based time series models have also sought to explicitly model error correlation structures to enhance forecasting performance.
@@ 4 Our Method
Models based on Gaussian likelihood are usually built on the assumption that $\boldsymbol{\eta}_{t}$ are independently distributed and following a multivariate Gaussian distribution. The log-likelihood of the distribution serves as the loss function for optimizing the model:
$\mathcal{L}=\sum_{t=1}^{T}\log p\left(\mathbf{z}_{t}\mid\theta\left(\mathbf{h}_{t}\right)\right)\propto\sum_{t=1}^{T}-\frac{1}{2}[\ln\lvert\boldsymbol{\Sigma}_{t}\rvert+\boldsymbol{\eta}_{t}^{\top}\boldsymbol{\Sigma}_{t}^{-1}\boldsymbol{\eta}_{t}].$ (6)
@@ 4.1 Training with Autocorrelated Errors
Building upon the concept of constructing mini-batches to address autocorrelated errors (Zheng et al., 2023a), we extend this approach to the multivariate scenario. In many existing deep probabilistic time series models, including GPVar (Salinas et al., 2019), each training batch encompasses a set of time series slices covering a temporal length of $P+Q$, where $P$ represents the conditioning range, and $Q$ denotes the prediction range. Gaussian likelihood is assessed individually at each time step within the prediction range. However, this simple approach does not account for the serial correlation of errors between successive time steps. To address this limitation, we organize $D$ consecutive batches into a mini-batch, with each batch covering a temporal length of $P+1$ (i.e., $Q=1$). In simpler terms, a traditional training batch comprises data $\mathbf{z}_{{T-P+1}:{T+Q}}$, while a mini-batch consists of data $\{\mathbf{z}_{{T-P+t}:{T+t}}\}_{t=1}^{D}$, where $T$ denotes a randomly sampled time step that separates the conditioning range and prediction range. It is evident that the mini-batch serves as a reconstruction of the conventional training batch, covering the same time horizon if $D=Q$. An example of the collection of target time series vectors in a mini-batch of size $D$ (the time horizon we use to capture serial correlation) is given by:
@@
4.2 Multistep-ahead Rolling Prediction
Autoregressive models conduct multistep-ahead forecasting iteratively. Specifically, the model generates a sample at each time step during prediction, serving as the input for the subsequent step. This process continues until the desired prediction range is reached. Our approach enhances this process in a similar way as in Zheng et al. (2023a) by offering additional calibration based on the learned correlation matrix $\boldsymbol{C}_{t}$. Assuming observations are available up to time step $t$ and considering that the errors collectively follow a multivariate Gaussian distribution in a mini-batch, the conditional distribution of $\boldsymbol{\eta}_{t+1}$ given errors in the past $(D-1)$ steps, can be derived as follows:
@@ 4.2 Multistep-ahead Rolling Prediction
where $\boldsymbol{\eta}_{\text{obs}}=\operatorname{vec}\left(\left[\boldsymbol{\eta}_{t-D+2},\ldots,\boldsymbol{\eta}_{t-1},\boldsymbol{\eta}_{t}\right]\right)\in\mathbb{R}^{(D-1)B}$ represents the set of residuals, accessible at forecasting step $t+1$. Here, $\boldsymbol{\Sigma}_{\text{obs}}$ is a $(D-1)B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{\text{obs}}$, and $\boldsymbol{\Sigma}_{*}$ is a $B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{t+1}$ and $\boldsymbol{\eta}_{\text{obs}}$, i.e., $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}=\begin{bmatrix}\boldsymbol{\Sigma}_{\text{obs}}&\boldsymbol{\Sigma}_{*}^{\top}\\ \boldsymbol{\Sigma}_{*}&\boldsymbol{\Sigma}_{t+1}\end{bmatrix}$. For conciseness, we omit the time index $t$ in $\boldsymbol{\Sigma}_{\text{obs}}$, $\boldsymbol{\Sigma}_{*}$ and $\boldsymbol{\eta}_{\text{obs}}$. Given that $\boldsymbol{\mu}_{t+1}$ is a deterministic output from the base model, a sample of the target variables $\tilde{\mathbf{z}}_{t+1}$ can be derived by initially drawing a sample $\tilde{\boldsymbol{\eta}}_{t+1}$ from Eq. (13) and then combining it with the predicted mean vector $\boldsymbol{\mu}_{t+1}$ by
@@
It is apparent that the adjusted distribution for ${\mathbf{z}}_{t+1}$ conditioned on the observed residuals $\boldsymbol{\eta}_{\text{obs}}$ is
@@
By considering the sample $\tilde{\boldsymbol{\eta}}_{t+1}$ as an observed residual, we can iteratively execute the process described in Eq. (13) to derive a trajectory of $\{\tilde{\mathbf{z}}_{t+q}\}_{q=1}^{Q}$. By repeating this procedure, we can generate samples representing the predictive distribution at each time step.